OpenStudy (anonymous):

Identify the # that needs to be added to both sides of the equation x(squared)-8x=-15

OpenStudy (anonymous):

to complete the square

OpenStudy (gw2011):

x^2-8x=-15 Add +15 to both sides and you get: x^2-8x+15=0 (x-5)(x-3)=0 x-5=0 x=5 x-3=0 x=3 Therefore, x=3 and x=5